OpenStudy (anonymous):

Which of the following are solutions to the equation below? Check all that apply. 5x2 + 3x - 2 = 0

OpenStudy (wolfboy):

Whats the fallowing?

OpenStudy (anonymous):

2/5

OpenStudy (anonymous):

-1

OpenStudy (lgbasallote):

i dont get what your comments say

OpenStudy (cwrw238):

ax^2 + bx + c =0 formula for roots is [-b +- sqrt(b^2 - 4ac)] / 2a plug values of a b and c and see what you get there will be two roots

OpenStudy (campbell_st):

the quadratic can be factorised (5x - 2)(x + 1)= 0 so x = 2/5 and -1 are the solutions